# Universal Anomaly Detection at the LHC: Transforming Optimal Classifiers and the DDD Method

### v3

@@ Abstract
We compare the performance of the DDD method with the Deep Robust One-Class Classification method (DROCC), which incorporates signals in the training process, and the Deep Support Vector Data Description (DeepSVDD) method, a well established and well performing method for anomaly detection. Results show that the effectiveness of each model varies by signal and channel, with DDD proving to be a very effective anomaly detector. We recommend the combined use of DeepSVDD and DDD for purely unsupervised applications, with the addition of flow models for improved performance when resources allow.
@@ 1 Introduction
Assuming a signal and supervised event classifiers: The search for these effects is like looking for a needle in a haystack. The production of new particles generally have a small cross-section and have to be filtered out of an enormous number of background events. Classification algorithms have been used for this purpose since the 1980s or 1990s, see e.g. Peterson:1993nk ; Stimpfl-Abele:1993xgr ; Ametller:1996ri ; Denby:1987rk and since 2014 Baldi:2014kfa , more and more deep learning algorithms have been used. Classification algorithms such as neural networks learn with labelled data (i.e. supervised) to distinguish signals and background events given physical features (such as the 4-vectors and object types) based on training data by adjusting weights to minimize a so-called loss function. Typically, the loss function is a binary cross entropy, that measures the difference of the true class probability distribution to the predicted class probability distribution. Minimizing this penalty function then leads to the classifier learning the ratio of probability distributions in the feature space between signal and background. In recent years, deep learning architectures have been shown to provide better classifiers than the boosted decision trees widely used in high-energy physics. More recently, fully connected deep learning architectures have been overtaken by graph networks Shlomi:2020gdn and even attention-based transformer architectures qu2024particle , inspired by their success in language models DBLP:journals/corr/VaswaniSPUJGKP17 .

### v4

@@ Abstract
We compare the performance of the DDD method with the Deep Robust One-Class Classification method (DROCC), which incorporates signals in the training process, and the Deep Support Vector Data Description (DeepSVDD) method, a well-established and well-performing method for anomaly detection. Results show that the effectiveness of each model varies by signal and channel, with DDD proving to be a very effective anomaly detector. We recommend the combined use of DeepSVDD and DDD for purely unsupervised applications, with the addition of flow models for improved performance when resources allow.
@@ 1 Introduction
Assuming a signal and supervised event classifiers: The search for these effects is like looking for a needle in a haystack. The production of new particles generally have a small cross-section and have to be filtered out of an enormous number of background events. Classification algorithms have been used for this purpose since the 1980s or 1990s, see e.g. Peterson:1993nk ; Stimpfl-Abele:1993xgr ; Ametller:1996ri ; Denby:1987rk and since 2014 Baldi:2014kfa , more and more deep learning algorithms have been used. Classification algorithms such as neural networks learn with labelled data (i.e. supervised) to distinguish signals and background events given physical features (such as the 4-vectors and object types) based on training data by adjusting weights to minimize a so-called loss function. Typically, the loss function is a binary cross entropy, that measures the difference of the true class probability distribution to the predicted class probability distribution. Minimizing this penalty function then leads to the classifier learning the ratio of probability distributions in the feature space between signal and background. In recent years, deep learning architectures have been shown to provide better classifiers than the boosted decision trees widely used in high-energy physics. More recently, fully connected deep learning architectures have been overtaken by graph networks Shlomi_2021 and even attention-based transformer architectures qu2024particle , inspired by their success in language models DBLP:journals/corr/VaswaniSPUJGKP17 .
@@ 5.2 Performance for the 4-top channel
In addition, the performance of the anomaly detection method, namely the DDD ParT+SM model, was evaluated under different jet energy scale (JES) variations applied to the 4-top signal ATLAS:2014hvo ; ATLAS:2017bje . The results, presented in Appendix B, indicate that scaling the jet energies up or down introduced only minor variations in the anomaly detection metrics. This suggests that the model does not mistake systematic JES shifts for true anomalies and remains robust under realistic energy scale uncertainties. This study confirms that the JES variations applied to the 4-top signal do not mimic the characteristics of the generic anomaly the model is designed to detect.
@@ A.1 Dark Machines channels
Appendix B JES studies
To investigate the potential impact of jet energy scale (JES) variations on the performance of the DDD ParT+SM model, a study was conducted in which the jet energy of signal (4-top) events was systematically scaled up and down by different percentages. The aim was to examine whether these systematic variations in jet energy could significantly affect anomaly detection results.
Fixed scaling factors of ±$\pm 2\%$, ±$\pm 3\%$, ±$\pm 5\%$, ±$\pm 8\%$, and ±$\pm 10\%$ were applied to the jet and b-jet energies, with the transverse momentum ($p_{T}$) recalculated accordingly.
The results, presented in Fig. 6, demonstrate that while scaling jet energies up or down introduces minor systematic changes in AUC and signal efficiency ($\varepsilon_{S}$), these variations remain relatively small and gradual across all tested scale percentages. This suggests that the model remains stable under reasonable JES variations.
Figure 6: Performance of the DDD ParT+SM in the 4-top channel with JES variations of up to 10%. The left plot shows AUC values, while the right plot displays signal efficiency ($\epsilon_{S}$) for a fixed background efficiency of 10%. The markers represent different scale types: ‘Up’ (blue) and ‘Down’ (red). The green marker corresponds to the original model without any applied scaling.
In Section 5.2, the results of this study are discussed in the context of assessing the robustness of the DDD method when systematic uncertainties, such as JES, are considered.